The heat of vaporization of methanol, CH3OH, is 35.20 kJ/mol. Its boiling




point is 64.6°C. What is the change in entropy for the vaporization of




methanol?

1
Expert's answer
2022-05-30T17:02:03-0400

Step 1 : given data

– standard enthalpy of vaporization (∆H°vap) = 35,2 kJ/mol;

– normal boling point (Tb) = 64,6°C

Step 2 : Convert Tb to Kelvin

We will use the following expression.

K = °C + 273

K = 64,6 + 273 = 337,6 K

Step 3 : Calculate the standard change entropy for the vaporization of methanol (∆S°vap)

The vaporization is the best phase transition from the liquid phase to vapor. We can calculate the standard change in entropy for the vaporization using the following expression.

∆S°vap = ∆H°vap / Tb

∆S°vap = (35,2 * 103 J/mol) / 337,6 K

∆S°vap = 104,2 J/mol•K ≈ 104 J/mol•K

ANSWER: 104 J/mol•K